Question 5 The comparative statements of Cullumber Company are presented here. CULLUMBER COMPANY ncome StatementS For the Years Ended December 31 2020 2019 Net sales Cost of goods sold Gross profit Selling and administrative expenses Income from operations Other expenses and losses $1,896,240 $1,756,200 1,011,700 744,500 484,700 259,800 1,064,240 832,000 505,700 326,300 Interest expense Income before income taxes Income tax expense Net income 23,300 303,000 93,300 $209,700 21,300 238,500 74,300 $164,200Balance Sheets December 31 Assets 2020 2019 Current assets Cash Debt investments (short-term) Accounts receivable Inventory $64,200 50,000 108,500 116,800 339,500 532,300 $1,045,900 $871,800 $60,100 74,000 123,500 127,300 384,900 661,000 Total current assets Plant assets (net) Total assets Liabilities and Stockholders Equity Current liabilities Accounts payable Income taxes payable $165,700 $151,100 43,300 194,400 212,000 406,400 44,800 210,500 232,000 442,500 Total current liabilities Bonds payable Total liabilities Stockholders equity Common stock ($5 par) Retained earnings 300,000 165,400 465,400 $1,045,900$871,800 290,000 313,400 603,400 Total stockholders equity Total liabilities and stockholders equityAll sales were on account. Net cash provided by operating activities for 2020 was $238,000. Capital expenditures were $135,000, and cash dividends were $61,700 Compute the following ratios for 2020, (Round free cash flow to 0 decimal places, eg. 5,275 and all other answers to 2 decimal places, e.g. 1.83 or 1.83%. Use 365 days for calculation.) (a) Earnings per share (b) Return on common stockholders equity (c) Return on assets (d) Current ratio (e) Accounts receivable turnover (f) Average collection period (g) Inventory turnover (h) Days in inventory (i) Times interest earned i) Asset turnover (k) Debt to assets ratio (I) Free cash flovw 3.5 240.89 % 1.82 1 252.83 times 1.44 days 202.71 times 1.80 days 14 times 361.18 times 42.30 % 41,30
Answer #1

Return on Common Stockholder's equity = Net income/Average common Stockholder's equity

Average common Stockholder's equity = (290000+300000+313400+165400/2) = 534400

Return on Common Stockholder's equity = 209700/534400 = 39.24%